Letsgo13:

You have a rectangular space where you plan to create an obstacle course for an animal. The area of the rectangular space is represented by the expression 6x2 − 16x. The width of the rectangular space is represented by the expression 2x.

jhonyy9:

so i think you ve missed details ... ???

KyledaGreat:

To find the length of the rectangular space, we can use the formula for the area of a rectangle: A = l*w where A is the area, l is the length, and w is the width. We are given that the area is 6x^2 - 16x, and the width is 2x. So we can substitute these values into the formula: 6x^2 - 16x = l * 2x Simplifying, we get: 3x - 8 = l So the length of the rectangular space is 3x - 8.